DIFFERENTIAL GEOMETRY/PDE SEMINAR

Tuesday, February 10, 2004 Thomson 135 1:30 pm (Note: special day and room!)

Hidden symmetry

Benson Farb (University of Chicago)

In this talk I will report on joint work with Shmuel Weinberger, addressing the question: which metrics on a Riemannian manifold have the most symmetry? Our main theorem is essentially a classification (up to finite index subgroups ) of the isometry groups of all closed Riemannian manifolds and their universal covers. I will try to explain this theorem, how we discovered it, and some of its many consequences. This talk should be acecssible to first year graduate students.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.